OpenStudy (anonymous):

P1 P2 M (3, 5) (-2, 0)

OpenStudy (anonymous):

@timo86m can u help me?

OpenStudy (anonymous):

.5,2.5

OpenStudy (anonymous):

how would i find this Find the coordinates of the center and the length of the radius of the following circle. (x + 3)2 + (y - 6)2 = 24 center: radius: <----

OpenStudy (anonymous):

use equation of circle and it will show -3,6 is center readious is sqrt(24)

OpenStudy (anonymous):

thanks!!!:D

OpenStudy (anonymous):

what would the center of this be? Find the coordinates of the center and the length of the radius of the following circle. (x + 2)2 + y2 = 10 center= 2,1??

OpenStudy (anonymous):

@TheViper If you see this i cant message you cuzz you only accept from people you fan but to make ze heart press alt+3 the 3 has to be on your numpad

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

(x + 2)2 + y2 = 10 center= 2,1?? -2,0 altho i think it will be zero :) radious sqrt(10)

OpenStudy (theviper):

Thanx @timo86m now I m ur fan:)

OpenStudy (anonymous):

woot :D

OpenStudy (theviper):

press alt+3 in equation editor??

OpenStudy (anonymous):

:O thanks timo!!!!!!!!!!!:D

OpenStudy (theviper):

\[\Huge{\color{gold}{\star \star}{\color{red}{I \heartsuit \LaTeX}}}\]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

no just in normal text editior viper

OpenStudy (theviper):

k! thank u:)

OpenStudy (theviper):

\[\Huge{\color{blue}{♥}\color{red}{♥}Thanx}\]

OpenStudy (theviper):

@timo86m

OpenStudy (anonymous):

:D aww